OpenStudy (anonymous):

Find out how long it takes a $2800.00 investment to double if it is invested at 9% compounded quarterly. Round to the nearest tenth of a year.

OpenStudy (anonymous):

this the question dude right

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

is there like answer choices right bro

OpenStudy (anonymous):

7.8, 8.2, 7.6, and 8 years

OpenStudy (anonymous):

2=(1.0225)^(4t) ln2=4t*ln1.0225 t=7.8 this one example

OpenStudy (anonymous):

7.8

OpenStudy (anonymous):

ok man

OpenStudy (anonymous):

is it right ?

OpenStudy (anonymous):

ce ya